# Very Massive Stars and High N/O: A Tale of the Nitrogen-enriched Super Star Cluster in the Sunburst Arc

### v1

@@ Abstract
The lensed Sunburst Arc ($z=2.369$) hosts a young ($\sim 2$$4\,\rm Myr$), massive ($M_{\star}\sim 10^{7}\,M_{\odot}$), compact ($R_{\rm eff}\sim 8\,\rm pc$) Lyman-continuum (LyC) leaking super star cluster, which powers a compact ($10\,\rm pc$), high-pressure nebula at sub-solar metallicity $\sim 0.2\,Z_{\odot}$ and with an anomalously elevated nitrogen-to-oxygen ratio $\log({\rm N/O})\sim-0.2$. We present semi-analytic models and 3D magnetohydrodynamic simulations with radiative feedback in an attempt to reproduce this system. The results indicate that the progenitor giant molecular cloud (GMC) may have $M_{\rm cloud}\gtrsim 3\times 10^{7}\,M_{\odot}$ and $R_{\rm cloud}\sim 70\,\rm pc$, corresponding to a surface density $\sim 10^{3}$$10^{4}\,M_{\odot}\,{\rm pc}^{-2}$. Incorporating feedback from individual Very Massive Stars (VMSs; $\geq 100\,M_{\odot}$) sampled from the Kroupa initial mass function, we find that their winds rapidly enrich $\sim 10^{4}\,M_{\odot}$ of nearby gas with nitrogen ($\sim 1\,$dex) and helium ($\sim 0.1$0.20.2\,dex). In the first 1133\,Myr, some cold gas falls to the system center where a central cluster builds up from sub-cluster mergers. There, the gas is photoionized, pressurized, and chemically enriched by the newly formed VMSs, before being radiatively expelled in the next $\sim 1\,\rm Myr$. We find that both VMS feedback and a high-surface-density progenitor GMC are necessary to reproduce the observed nebular properties, such as high N/O, high pressure, and stellar proximity. Low metallicity ($Z\leq 0.004$) may be essential to avoid overproduction of carbon from WC stars. Such enrichment processes localized to compact starburst events may have caused strong nitrogen emission from dense ionized gas as observed in high-redshift galaxies such as GN-z11 and GS_3073.
@@
software: Gizmo (Hopkins, 2015)
1 Introduction
@@
Strong gravitational lensing enables a magnified view of distant galaxies. The Sunburst Arc at $z\approx 2.369$ is one of the brightest lensed galaxies (Dahle et al., 2016). Inside that galaxy is a compact star-forming clump that features a high Lyman continuum (LyC) escape fraction (Rivera-Thorsen et al., 2017, 2019), which is determined to be a $\sim 2$$4\,{\rm Myr}$-old (Chisholm et al., 2019; Pascale et al., 2023; Meštrić et al., 2023; Rivera-Thorsen et al., 2024) super star cluster with a stellar mass in excess of $\sim 10^{7}\,M_{\odot}$ (Pascale et al., 2023; Rivera-Thorsen et al., 2024) and an effective radius of $\sim 8\,\rm pc$ (Vanzella et al., 2022; Meštrić et al., 2023). The gas metallicity is $Z\sim 0.2\,Z_{\odot}$ (Pascale et al., 2023; Rivera-Thorsen et al., 2024; Welch et al., 2025). Moreover, Pascale et al. (2023) found through nebular spectroscopy highly-pressurized ($p/k_{\rm B}\gtrsim 10^{9}\,\rm K\,cm^{-3}$) photo-ionized gas within $\sim 10\,$pc of the cluster center, which features a high nitrogen abundance $\log(\rm N/O)\sim-0.21$, while the outer, lower-pressure gas is not enriched. This is in strong excess of the median ISM N/O value for its metallicity according to empirical relations (Nicholls et al., 2017).
@@
Similar nitrogen abundance anomalies have been observed from systems at various cosmic redshifts (Marques-Chaves et al., 2024). Remarkably, several high-redshift galaxies (especially GN-z11 and GS_3073, e.g., Bunker et al., 2023; Cameron et al., 2023; Ji et al., 2024, and more references in § 5) and AGNs (Isobe et al., 2025) show nitrogen-enriched nebulae in JWST observations. Compared with the Sunburst Arc super star cluster, the nitrogen enhancement for some of these galaxies is more extreme ($\log(\rm N/O)\gtrsim 0.5$, e.g., Ji et al., 2024). Even in the Sunburst Arc, another compact young star cluster is found to host an N-enriched, ultra-dense ($\log n_{\rm e}=7$88) nebula (Pascale & Dai, 2024) (also see Choe et al., 2025). Toward low redshifts, nitrogen emitters appear exceedingly rare (Bhattacharya & Kobayashi, 2025), but the central young starburst of the blue compact dwarf galaxy Mrk 996 is one local example with a very dense enriched nebula (James et al., 2009). The dwarf galaxy NGC 5253 hosts another $\sim 10^{5}\,M_{\odot}$ embedded super star cluster with localized nitrogen enrichment (Smith et al., 2016) (also see Abril-Melgarejo et al., 2024; Pruijt et al., 2025).
@@
Several stellar sources of pollution have been proposed to explain the nitrogen enrichment, including Wolf-Rayet (WR) stars (Kobayashi & Ferrara, 2024; Flury et al., 2025), temporally differential winds that are enriched by core-collapse supernovae (Rizzuti et al., 2025) or asymptotic giant branch (AGB) stars (D’Antona et al., 2023; McClymont et al., 2025; Bhattacharya & Kobayashi, 2025), OB stars (Tapia et al., 2024), Very Massive Stars (VMSs; Vink, 2023), and Supermassive Stars (Charbonnel et al., 2023; Nandal et al., 2025; Gieles et al., 2025). An intriguing observed fact is that these nitrogen enriched nebulae always exhibit unusually high electron density $\log n_{\rm e}\gtrsim 5$ (James et al., 2009; Pascale et al., 2023; Ji et al., 2024; Topping et al., 2024; Yanagisawa et al., 2024; Isobe et al., 2025), which hints at gaseous environments subject to strong radiative or hydrodynamic compression in the vicinity of young massive stars (Pascale et al., 2023).
@@
Given the potential connections between high-redshift nitrogen-enriched galaxies and globular cluster stars with similar abundance anomalies (Isobe et al., 2023; Ji et al., 2025), understanding the LyC cluster of the Sunburst Arc at Cosmic Noon can shed crucial light on the very high-$z$ nitrogen-enriched galaxies. In particular, the LyC cluster is $\gtrsim 100$ times more massive than known young massive clusters of the Milky Way and the Magellenic Clouds. For its young age, a rapid mechanism of enrichment is required (Pascale et al., 2023; Rivera-Thorsen et al., 2024).
@@
In this study, we focus on the scenario associated with VMSs (Vink, 2023). VMSs are massive ($\gtrsim 100\,M_{\odot}$) and short-lived ($2-3\,\rm Myr$) stars (Vink et al., 2015) that are beyond the normally assumed mass range of an initial mass function (IMF) (e.g., Kroupa, 2002). Their abundance may be enhanced with a top-heavy IMF, which is theoretically expected in the dense star-forming environments of globular cluster progenitors (Matzner, 2024). Unlike less massive stars, VMSs can shed order unity of their initial mass through winds on the main sequence (Vink et al., 2011; Vink, 2018; Sabhahit et al., 2022). Moreover, they are typically rotating and highly convective, making the surface abundance of elements different from that of less massive stars (Higgins et al., 2023, 2024). In fact, they are often spectroscopically classified as WNh stars for prominent emission features from their strong winds. As such, VMS winds can be a plausible source of element enrichment for the early phase of chemical evolution of the ISM (Lahén et al., 2024). Given the young age of the Sunburst Arc LyC cluster, VMSs offer a rapid route to reach the high nitrogen abundance (Vink, 2023).
@@
Despite a number of previous works that explain the nitrogen enrichment of the Sunburst Arc LyC cluster, direct simulations on cluster formation, gas dynamical evolution, and chemical self-enrichment are required to reveal details of this process. In this study, we follow the method as presented in Shi & Murray (2025) to run simulations that integrate FIRE-3 star-formation physics (Hopkins et al., 2014, 2018a, 2023) and the evolution of individual VMSs, allowing us to explicitly model VMS feedback. Here, we further implement stellar evolution tracks from the PARSEC dataset (Nguyen et al., 2022; Costa et al., 2025) such that chemical yields from VMSs are included in the simulation.
This article is organized as follows. As a proof of concept, in § 2, we first compile the PARSEC data to demonstrate that VMSs can explain the rapid nitrogen enrichment in young star clusters like the Sunburst Arc LyC cluster. We then introduce our simulations, starting with the method in § 3, and then the results in § 4, where we present our fiducial runs that approximately reproduce the features of the LyC cluster, as well as control simulations to study the dynamics of star formation and the feedback from VMSs. In § 5, we also discuss the implications of this scenario on the high-redshift nitrogen-enriched galaxies. We finally conclude in § 6.
2 Chemical enrichment by stellar winds: a toy model
Most chemical elements heavier than helium (Alpher et al., 1948) are forged in stars (Burbidge et al., 1957). The chemical feedback from massive stars can be modeled in star formation simulations with the help of stellar evolution codes. In this article, we implement our sub-grid model based on the public evolution tracks, PARSEC v2.0 (Nguyen et al., 2022; Costa et al., 2025). The PARSEC dataset provides the full evolution history of non-rotating, intermediate to very massive stars for a range of initial metallicity $10^{-11}-0.03$ (Costa et al., 2025), largely covering the necessary parameter space required by our simulation.
@@ 2 Chemical enrichment by stellar winds: a toy model
Here $\langle m\rangle$ is the mean stellar mass obtained from the IMF, $\dot{m}$ is the mass loss rate of the star, and $X_{i}(t)$ is the surface abundance of element $i$. All these quantities are available from the PARSEC data.
@@
Simulations of star formation from giant molecular clouds (GMCs) suggest that the SFR peaks roughly at one free-fall time (Grudić et al., 2018; Shi & Murray, 2025). As a toy model, we assume that the free-fall time is $t_{\rm ff}=1\,\rm Myr$ and the SFR has a Gaussian with a standard deviation of $0.75\,\rm Myr$. We also assume a high star formation efficiency of $0.4$. The star formation history is initialized with $M_{\star}=0$ at $t=0$.
As star formation begins, the pristine gas is partially converted into stars and partially mixed with stellar ejecta. Star formation simulations also suggest that a high fraction of the gas is removed from the cluster due to stellar feedback (Grudić et al., 2018). Even for a very compact globular cluster progenitor (like the Sunburst Arc LyC cluster), radiative feedback is powerful enough that it can push the gas away while emitting LyC photons (Menon et al., 2024, 2025). This depletion of gas from the cluster is not instantaneous, and in our toy model, we assume the mass depletion rate is proportional to the SFR, but delayed by $1\,\rm Myr$. We also assume that depletion is very efficient, such that only 2% of the initial GMC mass remains in the cluster as gas at the late stage. This leads to the case where the mixed gas is largely dominated by the elemental abundances of the stellar winds, a conclusion also suggested by previous analytic works. For example, Tapia et al. (2024) found that $M_{\rm gas}/M_{\star}\sim 0.01$ at the time of N/O anomaly in their model, while our assumptions suggest $M_{\rm gas}/M_{\star}=0.02/0.4=0.05$, which is more conservative.
Fig. 1 shows the evolution of total stellar mass and the pristine gas mass (including two cases, with or without pristine-gas outflow). On top of these, we show the evolution of the stellar wind ejecta mass $M_{\rm ejecta}$ by assuming different combinations of the GMC’s initial metallicity $Z$ and the IMF’s high-mass end cutoff $m_{\rm cut}$ (to include/exclude VMSs), which are compiled with Eq. (2). For both $Z=0.014$ and $Z=0.004$, VMSs in the mass range $100,300]\,M_{\odot}$ boost the wind ejecta mass by a factor of $\sim 2$. Given the star formation efficiency $0.4$ in this toy model, we find $\sim 10\%$ ($\sim 3\%$) of the stellar mass will be converted to winds for $Z=0.014$ ($Z=0.004$), which is comparable to the left-over pristine gas with feedback-driven outflow that we assumed in the toy model.
@@
Figure 1: Mass evolution of star formation and feedback in the toy model (§ 2), including the total mass of stars $M_{\star}$ and the mass of the pristine gas $M_{\rm pristine}$ (without or with feedback-driven outflow). Moreover, we show the cumulated mass of wind ejecta $M_{\rm ejecta}$ from stellar populations of different initial metallicity $Z$ and IMF cutoff $m_{\rm max}$.
2.1 Abundances of pristine gas

### v2

@@ Abstract
The lensed Sunburst Arc ($z=2.369$) hosts a young ($\sim 2$$4\,\rm Myr$), massive ($M_{\star}\sim 10^{7}\,M_{\odot}$), compact ($R_{\rm eff}\sim 8\,\rm pc$) Lyman-continuum (LyC) leaking super star cluster, which powers a compact ($10\,\rm pc$), high-pressure nebula at sub-solar metallicity $\sim 0.2\,Z_{\odot}$ and with an anomalously elevated nitrogen-to-oxygen ratio $\log({\rm N/O})\sim-0.2$. We present semi-analytic models and 3D magnetohydrodynamic simulations in an attempt to reproduce this system. The results indicate that the progenitor giant molecular cloud (GMC) may have $M_{\rm cloud}\gtrsim 3\times 10^{7}\,M_{\odot}$ and $R_{\rm cloud}\sim 70\,\rm pc$, corresponding to a surface density $\sim 10^{3}$$10^{4}\,M_{\odot}\,{\rm pc}^{-2}$. Incorporating feedback from individual very massive stars (VMSs; $\geq 100\,M_{\odot}$) sampled from the Kroupa initial mass function, we find that their winds rapidly enrich $\sim 10^{4}\,M_{\odot}$ of nearby gas with nitrogen ($\sim 1\,$dex) and helium ($\sim 0.1$0.20.2\,dex). In the first 1133\,Myr, some cold gas falls to the system center where a central cluster builds up from sub-cluster mergers. There, the gas is photoionized, pressurized, and chemically enriched by the newly formed VMSs, before being radiatively expelled in the next $\sim 1\,\rm Myr$. We find that both VMS feedback and a high-surface-density progenitor GMC are necessary to reproduce the observed nebular properties, such as high N/O, high pressure, and stellar proximity. Low metallicity ($Z\leq 0.004$) may be essential to avoid overproduction of carbon from WC stars. Such enrichment processes localized to compact starbursts may have caused strong nitrogen emission from dense ionized gas as observed in high-redshift galaxies such as GN-z11 and GS_3073.
@@
software: Gizmo (Hopkins_2015MNRAS.450...53H)
I Introduction
@@
Strong gravitational lensing enables a magnified view of distant galaxies. The Sunburst Arc at $z\approx 2.369$ is one of the brightest lensed galaxies (DahleAghanimGuennou_2016A&A...590L...4D). Inside that galaxy is a compact star-forming clump that features a high Lyman continuum (LyC) escape fraction (Rivera-ThorsenDahleGronke_2017A&A...608L...4R; Rivera-ThorsenDahleChisholm_2019Sci...366..738R), which is determined to be a $\sim 2$$4\,{\rm Myr}$-old (ChisholmRigbyBayliss_2019ApJ...882..182C; PascaleDaiMcKee_2023ApJ...957...77P; MestricVanzellaUpadhyaya_2023A&A...673A..50M; Rivera-ThorsenChisholmWelch_2024A&A...690A.269R) super star cluster with a stellar mass in excess of $\sim 10^{7}\,M_{\odot}$ (PascaleDaiMcKee_2023ApJ...957...77P; Rivera-ThorsenChisholmWelch_2024A&A...690A.269R) and an effective radius of $\sim 8\,\rm pc$ (VanzellaCastellanoBergamini_2022A&A...659A...2V; MestricVanzellaUpadhyaya_2023A&A...673A..50M). The gas metallicity is $Z\sim 0.2\,Z_{\odot}$ (PascaleDaiMcKee_2023ApJ...957...77P; Rivera-ThorsenChisholmWelch_2024A&A...690A.269R; WelchRivera-ThorsenRigby_2025ApJ...980...33W). Moreover, PascaleDaiMcKee_2023ApJ...957...77P found through nebular spectroscopy highly-pressurized ($p/k_{\rm B}\gtrsim 10^{9}\,\rm K\,cm^{-3}$) photo-ionized gas within $\sim 10\,$pc of the cluster center, which features a high nitrogen abundance $\log(\rm N/O)\sim-0.21$, while the outer, lower-pressure gas is not enriched. This is in strong excess of the median ISM N/O value for its metallicity according to empirical relations (NichollsSutherlandDopita_2017MNRAS.466.4403N).
@@
Similar nitrogen abundance anomalies have been observed from systems at various cosmic redshifts (Marques-ChavesSchaererKuruvanthodi_2024A&A...681A..30M). Remarkably, several high-redshift galaxies (especially GN-z11 and GS_3073, e.g., BunkerSaxenaCameron_2023A&A...677A..88B; CameronKatzRey_2023MNRAS.523.3516C; JiUblerMaiolino_2024MNRAS.535..881J, and more references in § V) and AGNs (IsobeMaiolinoD'Eugenio_2025MNRAS.541L..71I) show nitrogen-enriched nebulae in JWST observations. Compared with the Sunburst Arc super star cluster, the nitrogen enhancement for some of these galaxies is more extreme ($\log(\rm N/O)\gtrsim 0.5$, e.g., JiUblerMaiolino_2024MNRAS.535..881J). Even in the Sunburst Arc, another compact young star cluster is found to host an N-enriched, ultra-dense ($\log n_{\rm e}=7$88) nebula (PascaleDai_2024ApJ...976..166P) (also see ChoeEmilRivera-ThorsenDahle_2025A&A...698A..16C). Toward low redshifts, nitrogen emitters appear exceedingly rare (BhattacharyaKobayashi_2025arXiv250811998B), but the central young starburst of the blue compact dwarf galaxy Mrk 996 is one local example with a very dense enriched nebula (JamesTsamisBarlow_2009MNRAS.398....2J). The dwarf galaxy NGC 5253 hosts another $\sim 10^{5}\,M_{\odot}$ embedded super star cluster with localized nitrogen enrichment (SmithCrowtherCalzetti_2016ApJ...823...38S) (also see Abril-MelgarejoJamesAloisi_2024ApJ...973..173A; PruijtMonreal-IberoWeilbacher_2025arXiv250907810P).
@@
Several stellar sources of pollution have been proposed to explain the nitrogen enrichment, including Wolf-Rayet (WR) stars (KobayashiFerrara_2024ApJ...962L...6K; FukushimaYajima_2024PASJ...76.1122F; FluryArellano-CordovaMoran_2025MNRAS.543.3367F), temporally differential winds that are enriched by core-collapse supernovae (RizzutiMatteucciMolaro_2025A&A...697A..96R) or asymptotic giant branch (AGB) stars (D'AntonaVesperiniCalura_2023A&A...680L..19D; McClymontTacchellaSmith_2025arXiv250708787M; BhattacharyaKobayashi_2025arXiv250811998B), OB stars (TapiaBekkiGroves_2024MNRAS.534.2086T), Very Massive Stars (VMSs; Vink_2023A&A...679L...9V), and Supermassive Stars (CharbonnelSchaererPrantzos_2023A&A...673L...7C; NandalWhalenLatif_2025ApJ...994L..11N; GielesPadoanCharbonnel_2025MNRAS.544..483G). An intriguing observed fact is that these nitrogen enriched nebulae always exhibit unusually high electron density $\log n_{\rm e}\gtrsim 5$ (JamesTsamisBarlow_2009MNRAS.398....2J; PascaleDaiMcKee_2023ApJ...957...77P; JiUblerMaiolino_2024MNRAS.535..881J; ToppingStarkSenchyna_2024MNRAS.529.3301T; YanagisawaOuchiWatanabe_2024ApJ...974..266Y; IsobeMaiolinoD'Eugenio_2025MNRAS.541L..71I), which hints at gaseous environments subject to strong radiative or hydrodynamic compression in the vicinity of young massive stars (PascaleDaiMcKee_2023ApJ...957...77P).
@@
Given the potential connections between high-redshift nitrogen-enriched galaxies and globular cluster stars with similar abundance anomalies (IsobeOuchiTominaga_2023ApJ...959..100I; JiBelokurovMaiolino_2025MNRAS.tmp.1982J), understanding the LyC cluster of the Sunburst Arc at Cosmic Noon can shed crucial light on the very high-$z$ nitrogen-enriched galaxies. In particular, the LyC cluster is $\gtrsim 100$ times more massive than known young massive clusters of the Milky Way and the Magellenic Clouds. For its young age, a rapid mechanism of enrichment is required (PascaleDaiMcKee_2023ApJ...957...77P; Rivera-ThorsenChisholmWelch_2024A&A...690A.269R).
@@
In this study, we focus on the scenario associated with VMSs (Vink_2023A&A...679L...9V). VMSs are massive ($\gtrsim 100\,M_{\odot}$) and short-lived ($2-3\,\rm Myr$) stars (VinkHegerKrumholz_2015HiA....16...51V) that are beyond the normally assumed mass range of an initial mass function (IMF) (e.g., Kroupa_2002Sci...295...82K). Their abundance may be enhanced with a top-heavy IMF, which is theoretically expected in the dense star-forming environments of globular cluster progenitors (Matzner_2024ApJ...975L..17M). Unlike less massive stars, VMSs can shed order unity of their initial mass through winds on the main sequence (VinkMuijresAnthonisse_2011A&A...531A.132V; Vink_2018A&A...615A.119V; SabhahitVinkHiggins_2022MNRAS.514.3736S). Moreover, they are typically rotating and highly convective, making the surface abundance of elements different from that of less massive stars (HigginsVinkHirschi_2023MNRAS.526..534H; HigginsVinkHirschi_2024MNRAS.533.1095H). In fact, they are often spectroscopically classified as WNh stars for prominent emission features from their strong winds. As such, VMS winds can be a plausible source of element enrichment for the early phase of chemical evolution of the ISM (LahenNaabSzecsi_2024MNRAS.530..645L). Given the young age of the Sunburst Arc LyC cluster, VMSs offer a rapid route to reach the high nitrogen abundance (Vink_2023A&A...679L...9V).
@@
Despite a number of previous works that explain the nitrogen enrichment of the Sunburst Arc LyC cluster, direct simulations on cluster formation, gas dynamical evolution, and chemical self-enrichment are required to reveal details of this process. In this study, we follow the method as presented in Shi_inprep to run simulations that integrate FIRE-3 star-formation physics (HopkinsKeresOnorbe_2014MNRAS.445..581H; HopkinsWetzelKeres_2018MNRAS.480..800H; HopkinsWetzelWheeler_2023MNRAS.519.3154H) and the evolution of individual VMSs, allowing us to explicitly model VMS feedback. Here, we further implement stellar evolution tracks from the PARSEC dataset (NguyenCostaGirardi_2022A&A...665A.126N; CostaShepherdBressan_2025A&A...694A.193C) such that chemical yields from VMSs are included in the simulation.
This article is organized as follows. In § II, we compile PARSEC data as a proof of concept to show that VMSs can account for the rapid nitrogen enrichment in young clusters such as the Sunburst Arc LyC cluster. We then describe our simulations: the methods in § III and the results in § IV, where we present fiducial runs reproducing the LyC cluster features and control runs exploring star formation and VMS feedback. Section V discusses the implications for high-redshift nitrogen-enriched galaxies, and Section VI summarizes our conclusions.
II Chemical enrichment by stellar winds: a toy model
Most chemical elements heavier than helium (AlpherBetheGamow_1948PhRv...73..803A) are forged in stars (BurbidgeBurbidgeFowler_1957RvMP...29..547B). In this article, we implement sub-grid stellar yield models based on the public evolution tracks, PARSEC v2.0 (NguyenCostaGirardi_2022A&A...665A.126N; CostaShepherdBressan_2025A&A...694A.193C), which provides the full evolution history of non-rotating, intermediate to VMSs for a range of initial metallicity $10^{-11}-0.03$ (CostaShepherdBressan_2025A&A...694A.193C), largely covering the necessary parameter space required by our simulation.
@@ II Chemical enrichment by stellar winds: a toy model
[…]
@@
Simulations of star formation from giant molecular clouds (GMCs) suggest that the SFR peaks roughly at one free-fall time (GrudicHopkinsFaucher-Giguere_2018MNRAS.475.3511G; Shi_inprep). As a toy model, we assume that the free-fall time is $t_{\rm ff}=1\,\rm Myr$ and the SFR has a Gaussian with a standard deviation of $0.75\,\rm Myr$. We also assume a high star formation efficiency of $0.4$. The star formation history is initialized with $M_{\star}=0$ at $t=0$.
As star formation proceeds, the pristine gas is partly converted into stars and partly mixed with stellar ejecta. Simulations indicate that stellar feedback removes a large fraction of gas from clusters (GrudicHopkinsFaucher-Giguere_2018MNRAS.475.3511G). Even in compact globular cluster progenitors such as the Sunburst Arc LyC cluster, radiative feedback can expel gas while producing LyC emission (MenonLancasterBurkhart_2024ApJ...967L..28M; MenonBurkhartSomerville_2025ApJ...987...12M). In our toy model, gas depletion follows the star formation rate with a $1\,\rm Myr$ delay and is assumed highly efficient, leaving only 2% of the initial GMC mass as gas at late times. Consequently, the mixed gas is dominated by stellar-wind ejecta, consistent with analytic estimates. For example, TapiaBekkiGroves_2024MNRAS.534.2086T obtained $M_{\rm gas}/M_{\star}\sim 0.01$ at the time of the N/O anomaly, while our assumption of $M_{\rm gas}/M_{\star}=0.02/0.4=0.05$ is slightly more conservative.
[…]
@@
Figure 1: Mass evolution of star formation and feedback in the toy model (§ II), including the total mass of stars $M_{\star}$ and the mass of the pristine gas $M_{\rm pristine}$ (without or with feedback-driven outflow). Moreover, we show the cumulated mass of wind ejecta $M_{\rm ejecta}$ from stellar populations of different initial metallicity $Z$ and IMF cutoff $m_{\rm max}$.
II.1 Abundances of pristine gas